solve the problem:
The worlds largest taco contained approximately 1 kg of onion for every 6.6 kg of grilled steak. The total weight of these two ingredients was 617.6 kg. To the nearest tenth of a kilogram, how many kilograms of each ingredient were used to make the taco?

I am struggling with turning the story problem into the actual equation.

Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

The ratio of onions to grilled steak was 1:6.6 kg.
.
In this case, 'k' is simply an constant multiplier.
.
1k + 6.6k = 617.6
.
7.6k = 617.6
.
k = 617.6/7.6
.
k = 81.3
.
So the taco contained about 81.3 kg of onions and 6.6*81.3 = 536.3 kg of grilled steak.
.
Done.

Answer by stanbon(75887)

The worlds largest taco contained approximately 1 kg of onion for every 6.6 kg of grilled steak. The total weight of these two ingredients was 617.6 kg. To the nearest tenth of a kilogram, how many kilograms of each ingredient were used to make the taco?
----------------
1 kg of onion for every 6.6 kg of grilled steak is
a ratio of 1:6.6
---
1:6.6 is the same as x:6.6x where x is the onion weight.
----
Equation:
x + 6.6x = 617.6
7.6x = 617.6
x = 81.26 kg (amount of onion in the taco)
---
6.6x = 6.6*81.26 = 536.32 kg (amount of grilled steak in the taco)
=================================================================
Cheers,
Stan H.
===========